At constant temperature 80% NO dissociatesinto N2 and O2 the equilibrium constant for 2NO N2+O2

Grade:12th pass

1 Answers

Vikas TU
7 years ago
             2NO ------------------->       N2       +       O2
t=0         a                                          0                  0
t=teq.     a-x                                       x/2               x/2
Now given, 80% gets dissociated then remaining NO is 20%
that is,
a – x = 0.2a
0.8a = x
Hence we get N2 and O2 as: => 0.4a each.
Hence Eqm. Constant =>  (0.4a)^2/(0.2a)^2 = > 2^2 = 4